OpenStudy (anonymous):

Find the general solution of the given nonhomogeneous differential equation: y''+2y'+y=e^(-x)*ln(x)

OpenStudy (anonymous):

Homogeneous part: \[y''+2y'+y=0~~\Rightarrow~~r^2+2r+1=0~~\Rightarrow~~r=-1\] So the homogeneous set of solution is \[y_c=C_1e^{-x}+C_2xe^{-x}\] For the nonhomogeneous part, use variation of parameters. You know two sample solutions, \(e^{-x}\) and \(xe^{-x}\).

OpenStudy (anonymous):

what is Yp? not sure how to replace ln(x) ?

OpenStudy (anonymous):

When you're using variation of parameters, you have \(y_{p_1}=e^{-x}\) and \(y_{p_2}=xe^{-x}\). First, check to make sure they're linearly independent: \[W(e^{-x},xe^{-x})=\begin{vmatrix}e^{-x}&xe^{-x}\\-e^{-x}&e^{-x}(1-x)\end{vmatrix}=e^{-2x}(1-x)+xe^{-2x}=e^{-2x}\not=0\] So yes, they're linearly independent.